5. A year ago, Fred purchased 300 shares of RPJ stock for $15 400. The stock is currently selling for $40 a share and Fred has decided to sell all of his shares. What is the rate of return that Fred has earned on this investment if he received a special dividend of $5 per share?
Answer #1

Purchase price = amount/shares = 15400/300=51.333

Return = ((selling price+dividend)/purchase price-1) = ((40+5)/51.333-1)=-12.34%